7] y – 3y – 4y = 15+t+ 3e4tpls use power series method
Answer #1


4t Solution given differential Equation -4_3y- 44 =15+t+30+ 0 let Power series solution y(t) = § anth of above diff. equatio o n o put n=2 band 3)(4)94 - 313) 93 - 492 – 3816 = 0 → 1294 - 993 – 492 = 24 94 = to [ 24 +993 + 492] 94 = to 124 + 9x4167+1